### v1

# DiffusionGPT[Uncaptioned image]: LLM-Driven Text-to-Image Generation System

Abstract
Diffusion models have opened up new avenues for the field of image generation, resulting in the proliferation of high-quality models shared on open-source platforms. However, a major challenge persists in current text-to-image systems are often unable to handle diverse inputs, or are limited to single model results. Current unified attempts often fall into two orthogonal aspects: i) parse Diverse Prompts in input stage; ii) activate expert model to output. To combine the best of both worlds, we propose DiffusionGPT, which leverages Large Language Models (LLM) to offer a unified generation system capable of seamlessly accommodating various types of prompts and integrating domain-expert models. DiffusionGPT constructs domain-specific Trees for various generative models based on prior knowledge. When provided with an input, the LLM parses the prompt and employs the Trees-of-Thought to guide the selection of an appropriate model, thereby relaxing input constraints and ensuring exceptional performance across diverse domains. Moreover, we introduce Advantage Databases, where the Tree-of-Thought is enriched with human feedback, aligning the model selection process with human preferences. Through extensive experiments and comparisons, we demonstrate the effectiveness of DiffusionGPT, showcasing its potential for pushing the boundaries of image synthesis in diverse domains.
Figure 1: We propose a unified generation system DiffusionGPT, which leverages Large Language Models (LLM) to seamlessly accommodating various types of prompts input and integrating domain-expert models for output. Our system is capable of parsing diverse forms of inputs, including Prompt-based, Instruction-based, Inspiration-based, and Hypothesis-based input types. It exhibits the ability to generate outputs of superior quality.
[…]
1 Introduction
Recent years have witnessed the prevalence of diffusion models[6] in image generation tasks, revolutionizing image editing, stylization, and other related tasks. Both DALLE-2[14] and Imagen[17] are extremely good at generating images from text prompts. However, their non-open source nature has hindered widespread popularization and corresponding ecological development. The first open-source text-to-image diffusion model, known as Stable Diffusion (SD)[16], which has rapidly gained popularity and widespread usage. Various techniques tailored for SD, such as Controlnet[27], Lora, further paved the way for the development of SD and foster its integration into various applications. SDXL[10] is the latest image generation model tailored to deliver exceptional photorealistic outputs with intricate details and artistic composition. Moreover, the impact of SD extends beyond technical aspects. Community platforms such as Civitai, WebUI, and LibLibAI have emerged as vibrant hubs for discussions and collaborations among designers and creators. The capability evolution of Stable Diffusion, the advancements in SD-customized technologies, and the thriving community platforms, has created an exciting and fertile ground for further advancements in image generation.
Despite making significant strides, current stable diffusion models face two key challenges when applied to realistic scenarios:
  • Model Limitation: While stable diffusion (such as SD1.5) demonstrate adaptability to a wide range of prompts, they exhibit poor performance in specific domains. Conversely, domain-specific models (such as SD1.5+Lora) excel in producing extreme generation performance within specific sub-fields but lack versatility.
  • Prompt Constraint: During the training of stable diffusion, text information typically consists of descriptive statements, such as captions. However, when utilizing these models, we encounter various prompt types, including instructions and inspirations. The current generation models struggle to achieve optimal generation performance for these diverse prompt types.
The existing mix-matches between stable diffusion models and realistic applications often manifest as limited results, poor generalization, and increased difficulty in their practical implementation. A series of research works have explored to address these challenges and bridge the gap. While SDXL has made notable advancements in improving specific-domain performance, achieving the ultimate performance in this regard remains elusive. Other approaches involve incorporating prompt engineering techniques or fixed prompt templates to enhance the quality of input prompts and improve the overall generation output. While these approaches have shown varying degrees of success in mitigating the aforementioned challenges, they do not provide a comprehensive solution. This leads us to pose a fundamental question: Can we create a unified framework to unleash prompt constraint and activate corresponding domain expert model ?
In order to address the aforementioned question, we propose DiffusionGPT, which leverages Large Language Model (LLM) to offer a one-for-all generation system that seamlessly integrates superior generative models and effectively parses diverse prompts. DiffusionGPT constructs a Tree-of-Thought (ToT) structure, which encompasses various generative models based on prior knowledge and human feedback. When presented with an input prompt, the LLM first parses the prompt and then guides the ToT to identify the most suitable model for generating the desired output. Furthermore, we introduce Advantage Databases, where the Tree-of-Thought is enriched with valuable human feedback, aligning the LLM’s model selection process with human preferences.
The contributions of this work can be summarized as:
  • New Insight: DiffusionGPT employs Large Language Model (LLM) to drive the entire text-to-image generation system. The LLM acts as the cognitive engine, processing diverse inputs and facilitating expert selection for output.
  • All-in-one System: DiffusionGPT provides a versatile and professional solution by being compatible with a wide range of diffusion models. Unlike existing approaches that are limited to descriptive prompts, our framework can handle various prompt types, expanding its applicability.
  • Efficiency and Pioneering: DiffusionGPT stands out for its training-free nature, allowing for easy integration as a plug-and-play solution. Through the incorporation of the Tree-of-Thought (ToT) and human feedback, our system achieves higher accuracy and pioneers a flexible process for aggregation of more experts.
  • High Effectiveness: DiffusionGPT outperforms traditional stable diffusion models, demonstrating significant advancements. By providing an all-in-one system, we offer a more efficient and effective pathway for community development in the field of image generation.
2 Related Work
2.1 Text-based Image Generation
Initially, Generative Adversarial Networks (GANs) [15, 26] were widely used as the primary approach for Text-based image generation. However, the landscape of image generation has evolved, and diffusion models [6] have emerged as a dominant framework, especially when integrated with text encoders such as CLIP [12] and T5 [13], enabling precise text-conditioned image generation. For instance, DAELL-2 [14] leverages CLIP’s image embeddings, derived from CLIP’s text embeddings through a prior model, to generate high-quality images. Similarly, Stable Diffusion [16] directly generates images from CLIP’s text embeddings. Imagen [17], on the other hand, utilizes a powerful language model like T5 [13] to encode text prompts, resulting in accurate image generation. Transformer-based architectures have also demonstrated their efficacy in generating images from textual inputs. CogView2 [4] and Muse [2] are notable examples of such models. To align text-to-image diffusion models with human preferences, recent methods [8, 1, 24] propose training diffusion models with reward signals. This ensures that the generated images not only meet quality benchmarks but also closely align with human intent and preferences. These cutting-edge approaches aim to improve the fidelity and relevance of generated images to better satisfy user requirements.
Figure 2: Overview of DiffusionGPT. The workflow of DiffusionGPT consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection, and Execution Generation. The four steps are shown from left to right and interact with LLM continuously. The upper side shows the detailed process of each step. The lower side shows the example of the whole workflow.
2.2 Large Language Models (LLMs) for Vision-Language Tasks
The field of natural language processing (NLP) has witnessed a significant transformation with the emergence of large language models (LLMs) [9, 3, 21], which have demonstrated remarkable proficiency in human interaction through conversational interfaces. To further enhance the capabilities of LLMs, the Chain-of-Thought (CoT) framework [22, 28, 7, 29] has been introduced. This framework guides LLMs to generate answers step-by-step, aiming for superior final answers. Recent research has explored innovative approaches by integrating external tools or models with LLMs [11, 18, 20, 23, 19]. For example, Toolformer [18] empowers LLMs with the ability to access external tools through API tags. Visual ChatGPT [23] and HuggingGPT [19] extend the capabilities of LLMs by enabling them to leverage other models to handle complex tasks that go beyond language boundaries. Similarly, Visual Programming [5] and ViperGPT [20] harness the potential of LLMs in processing visual objects by utilizing programming languages to parse visual queries. Drawing inspiration from these endeavors, we embrace the concept of LLMs as versatile tools and leverage this paradigm to guide T2I models to generate high-quality images.
3 Methodology
Figure 3: Details of prompts during interactions with the ChatGPT[9]. Before being inputted into the ChatGPT, the slots “{}” in figure are uniformly replaced with the corresponding text values.
DiffusionGPT is an all-in-one system specifically designed to generate high-quality images for diverse input prompts. Its primary objective is to parse the input prompt and identify the generative model that produces the most optimal results, which is high-generalization, high-utility, and convenient.
DiffusionGPT composes of a large language model (LLM) and various domain-expert generative models from the open-source communities (e.g. Hugging Face, Civitai). The LLM assumes the role of the core controller and maintains the whole workflow of the system, which consists of four steps: Prompt Parse, Tree-of-thought of Models of Building and Searching, Model Selection with Human Feedback, and Execution of Generation. The overall pipeline of DiffusionGPT is shown in Figure 2.
3.1 Prompt Parse
The Prompt Parse Agent plays a pivotal role in our methodology as it utilizes the large language model (LLM) to analyze and extract the salient textual information from the input prompt. Accurate parsing of the prompt is crucial for effectively generating the desired content, given the inherent complexity of user input. This agent is applicable to various types of prompts, including prompt-based, instruction-based, inspiration-based, hypothesis-based, etc.
Prompt-based: The entire input is used as the prompt for generation. For example, if the input is “a dog” the prompt used for generation would be “a dog”.
Instruction-based: The core part of the instruction is extracted as the prompt for generation. For instance, if the input is “generate an image of a dog”, the recognized prompt would be “an image of a dog”.
Inspiration-based: The target subject of the desire is extracted and used as the prompt for generation (e.g., Input: “I want to see a beach”; Recognized: “a beach”).
Hypothesis-based: It involves extracting the hypothesis condition (“If xxx, I will xxx”) and the object of the forthcoming action as the prompt for generation. For instance, if the input is “If you give me a toy, I will laugh very happily”, the recognized prompt would be “a toy and a laugh face”.
By identifying these forms of prompts, Prompt Parse Agent allows DiffusionGPT to accurately recognize the core content that the user wants to generate, while mitigating the influence of noisy text. This process is crucial for selecting appropriate generative models and achieving high-quality generation results.
3.2 Tree-of-thought of Models
Following the prompt parsing stage, the subsequent step involves selecting appropriate generative models from an extensive model library to generate the desired images. However, considering the large number of models available, it is impractical to input all models simultaneously into the large language model (LLM) for selection. Additionally, as different models may exhibit similarities in their generation spaces, accurately identifying the most suitable model through a single fuzzy match across the entire model library becomes challenging. To address this issue and identify the optimal model, we propose the utilization of a model tree based on the concept of Tree-of-Thought (TOT).By leveraging the search capabilities of the model tree, we can narrow down the candidate set of models and enhance the accuracy of the model selection process.
Constructing the Model Tree using TOT. The Tree-of-Thought (TOT) of Model Building Agent is employed to automatically build the Model Tree based on the tag attributes of all models. By inputting the tag attributes of all models into the agent, it analyzes and summarizes potential categories derived from the Subject Domain and Style Domain. The Style categories are then incorporated as subcategories within the Subject category, establishing a two-layer hierarchical tree structure. Subsequently, all models are assigned to suitable leaf nodes based on their attributes, thereby completing the comprehensive model tree structure. The figure below illustrates the visual representation of the model tree. As the model tree is automatically constructed by the agent, this approach ensures convenient extensibility for incorporating new models. Whenever new models are added, agent seamlessly places them in the appropriate position within the model tree based on their attributes.
Searching the Model Tree using TOT. The search process within the Model Tree, based on the Tree-of-Thought (TOT) of Models Searching Agent, aims to identify a candidate set of models that closely align with the given prompt. This search method employs a breadth-first approach, systematically evaluating the best subcategory at each leaf node. At each level, the categories are compared against the input prompt to determine the category that exhibits the closest match. This iterative process continues to derive the candidate set for the subsequent leaf node, and the search progresses until reaching the final node, where the candidate set of models is obtained. This candidate set of models serves as the basis for model selection in the subsequent stage.
3.3 Model Selection
The model selection stage aims to identify the most suitable model for generating the desired image from the candidate set obtained in the previous stage. This candidate set represents a subset of the entire model library, consisting of models that exhibit relatively high matching degrees with the input prompt. However, the limited attribute information available from open-source communities poses challenges in precisely determining the best model while providing detailed model information to the large language model (LLM). To address this, we propose a Model Selection Agent that leverages human feedback and utilizes advantage database technology to align the model selection process with human preferences.
For the advantage database, we employ a reward model to calculate scores for all model-generated results based on a corpus of 10,000 prompts, storing the score information. Upon receiving an input prompt, we calculate the semantic similarity between the input prompt and the 10,000 prompts, identifying the top 5 prompts with the highest similarity. Subsequently, the Model Selection Agent retrieves the precomputed performance of each model for these prompts from an offline database and selects the top 5 models for each selected prompt. This process produces a candidate set of 5x5 models.
The agent then intersects the model set with the model candidate set obtained in the TOT of models stage, focusing on models with higher occurrence probabilities and relatively higher rankings. These models are ultimately chosen as the final selection for model generation.
3.4 Execution of Generation
Once the most suitable model has been selected, the chosen generative model is utilized to generate the desired images using the obtained core prompts.
Prompt Extension. To enhance the quality of prompts during the generation process, a Prompt Extension Agent is employed to augment the prompt. This agent leverages prompt examples from the selected model to automatically enrich the input prompt. The example prompts and the input prompts are both sent to LLM in the in-context learning paradigm. Specially, this agent incorporates rich descriptions and detailed vocabulary to the input prompts following the sentence pattern of example prompts. For example, if the input prompt is “an image of a laughing woman, fashion magazine cover” and example prompt is “fashion photography portrait of woman avatar, 1girl in blue lush Alien Rainforest with flowers and birds, fantasy, octane render, hdr, Dolby Vision, (intricate details, hyperdetailed:1.2), (natural skin texture, hyperrealism, soft light:1.2), fluffy short hair, sharp focus, night, necklace, Chinese mythology, cleavage, medium breasts, sci-fi headband, looking at viewer, best quality, perfect body”, the Prompt Extension Agent enhances it to a more detailed and expressive form such as: “The woman on the magazine cover is laughing joyfully, her eyes twinkling with delight. She is wearing a fashionable outfit that accentuates her curves, and her hair is styled in a way that complements her features”. This augmentation significantly improves the quality of the generated outputs.
4 Experiments
Figure 4: When comparing SD1.5-based DiffusionGPT with SD15[16], it is observed that DiffusionGPT excels in generating more realistic results at a fine-grained level for categories such as humans and scenes. The generated images demonstrate improved visual fidelity, capturing finer details and exhibiting a higher degree of realism compared to SD15.
4.1 Settings
In our experimental setup, the primary large language model (LLM) controller employed was ChatGPT[9], specifically utilizing the text-davinci-003 version, which is accessible through the OpenAI API. To facilitate the guidance of LLM responses, we adopted the LangChain framework, which effectively controlled and directed the generated outputs. For the generation models utilized in our experiments, we selected a diverse range of models sourced from the Civitai and Hugging Face communities. The selection process involved choosing the most popular models across different types or styles available on these platforms.
Table 1: Quantitative Results: We evaluate the aesthetic score with image-reward and aesthetic score. Compare to “SD15” and “Random” select expert models to output, both Tree-of-Thoughts and Human Feedback can improve the aesthetic of generated images and human alignment.
Method Image-reward Aes score
SD15 0.28 5.26
Random 0.45 5.50
DiffusionGPT wo HF 0.56 5.62
DiffusionGPT 0.63 5.70
4.2 Qualitative Results
4.2.1 Visualization of SD1.5 Version
To assess the efficacy of our system, we performed a comprehensive evaluation by comparing its generation performance against the baseline method, SD 1.5[16]. SD 1.5 serves as the foundational model for various specialized community models. The comparative results are presented in Figure 4. We conducted a detail analysis of four distinct prompt types and compared them along two key dimensions: semantic alignment and image aesthetics.
Figure 5: Comparison of SDXL version of DiffusionGPT with baseline SDXL[10]. All generated iamges are $\times×1024$ pixels.
Figure 6: Comparison of DiffusionGPT-Xl with base model.
Figure 7: User Study: Comparing DiffusionGPT with SD1.5. Users strongly prefer expert models selected by DiffusionGPT over the baseline in terms of prompts from all 10 categories.
Upon careful examination of the results, we have identified two notable issues with the base model: i) Semantic Lack: The base model’s generated images exhibit a limited focus on specific semantic classes derived from the input prompt, resulting in an incomplete capture of the overall semantic information. This limitation is particularly evident in all kinds of prompt types, where the base model struggles to effectively generate objects related to “man, chef, children, and toyland”. ii) Poor performance on human-related targets: The base model faces challenges in generating accurate facial and body details for human-related objects, leading to subpar image quality. This deficiency becomes apparent when comparing the aesthetic qualities of images depicting “girl and couple”.
In contrast, DiffusionGPT effectively addresses these limitations. The images produced by our system demonstrate relatively complete representations of the target regions, successfully capturing the semantic information encompassing the entire input prompt. Examples such as “man who whistles tunes pianos” and “a snowy wonderland where children build snowmen and have snowball fights” showcase the capacity of our system to encompass the broader context. Furthermore, our system excels in generating more detailed and accurate image for human-related objects. This is exemplified by the prompt “a romantic couple sharing a tender moment under a starry sky”.
Figure 8: Ablation study of DiffusionGPT. The random selection is the baseline method for generating images. The TOT or TOT+HF represent the performance of different agents.
Figure 9: Ablation study of Prompt Extension. The extension aims to provide the riched prompts that produces higher quality images.
4.2.2 Visualization of SDXL Version.
With the advancements in publicly available universal generation models, the newly improved method SD XL[10] has emerged as a promising approach, demonstrating superior generation results. To further enhance our system, we have developed an upgraded version by integrating various open-source community models based on SD XL. In order to evaluate the performance of our system, we compared it against SD XL, as depicted in Figure 5. It is important to note that all output images have a resolution of 1024x1024, and four distinct types of prompts were generated for the purpose of comparison. Upon careful analysis, it becomes apparent that SD XL occasionally exhibits a partial loss of semantic information in specific cases. For instance, the generated results for prompts involving “3D tiger” in the prompt-based category or “flying cars” in the instruction-based category may lack accurate representation. In contrast, our system excels in producing more precise and visually appealing representations. Noteworthy examples include the generation of “a white towel with a cartoon cat” and the depiction of a “starry sky”.
4.3 Quantitative Results
The alignment between user preferences and the quantitative findings presented in Table 1 serves as strong evidence of the robustness and effectiveness of DiffusionGPT. To further evaluate the different generated results, we employed the aesthetic predictor and human feedback related reward model. By comparing the effects of our basic version with the baseline model SD1.5, the results in Table 1 demonstrate that our overall framework outperforms SD15 in terms of image-reward and aesthetic score, achieving improvements of 0.35% and 0.44% respectively.
4.4 Ablation Study
4.4.1 Tree-of-Thought and Human Feedback
To validate the effectiveness of the designed components, we conducted a visual analysis of the performance achieved by incorporating different modules, as illustrated in Figure 8. In the figure, the variant labeled as “Random” represents the random sampling model. Notably, the random selection model generates a considerable number of images that do not align with the input prompts and lack semantic coherence. However, as we gradually incorporate the tree-of-thought (TOT) and human feedback (HF) modules into our system, a significant improvement in the quality of the generated images becomes evident. The images generated with the inclusion of TOT and HF modules exhibit enhanced realism, semantic alignment with the input prompts, and higher aesthetic appeal. This visual analysis showcases the advantage of our system in selecting superior models through the integration of TOT and HF components.
4.4.2 Prompt extension
To evaluate the effectiveness of the prompt extension agent, we conducted a comparison between the generation results using the original prompt and the extended prompt as inputs, as depicted in Figure 9. The extended prompts were designed to provide a richer and more detailed description of the desired image. Upon analysis, we observed that the extended prompts yielded notable improvements in the aesthetics and level of detail of the generated images. The inclusion of additional information in the extended prompts allowed for the generation of more visually appealing and artistically enhanced images.
4.5 User Study
To obtain real human preferences for the generated images, we conducted a user study comparing our model against a baseline model. We utilized image captions sourced from PartiPrompts[25], randomly selecting 100 prompts and generating four images for each prompt. Subsequently, we collected feedback from 20 users, who were asked to rate the superiority or equality of the images. This process resulted in approximately 400 votes for each base model (SD15 and SD XL). As depicted in Figure 7 and Figure 6, the results of the user study consistently revealed a clear preference for our model over the baseline. The users consistently expressed a distinct preference for the images generated by our model, indicating that they perceived them to be of higher quality or superiority compared to the baseline.
4.6 Limitations and Further works
Although DiffusionGPT has demonstrated the ability to generate high-quality images, there are still several limitations and our future plans are in the following ways:
Feedback-Driven Optimization We aim to incorporate feedback directly into the optimization process of LLM, enabling more refined pompt parse and model selection.
Expansion of Model Candidates To further enrich the model generation space and achieve more impressive results, we will expand the repertoire of available models.
Beyond Text-to-Image Tasks We intend to apply our insight to a broader set of tasks, including controllable generation, style migration, attribute editing, etc.
5 Conclusion
We propose Diffusion-GPT, a one-for-all framework that seamlessly integrates superior generative models and efficiently parses diverse prompts. By leveraging Large Language Models (LLMs), Diffusion-GPT gains insights into the intent of input prompts and selects the most suitable model from a Tree-of-Thought (ToT) structure. This framework offers versatility and exceptional performance across different prompts and domains, while also incorporating human feedback through Advantage Databases. To sum up, Diffusion-GPT is training-free and can be easily integrated as a plug-and-play solution, offers an efficient and effective pathway for community development in this field.